OpenStudy (emsxx3_):

HELP FOR MEDAL AND FAN The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2). A. y = 5x + 13 B. y = negative 1 over 5x + 7 over 5 C. y = − 1 over 5x − 7 over 5 D. y = 5x − 17

OpenStudy (freckles):

Can you identify the slope of the linear equation y=5x+4?

OpenStudy (yanasidlinskiy):

The slope of this line is 5. The slope of a perpendicular line is the negative reciprocal: -1/5 Now use the point-slope formula to find the equation: y - y1 = m (x - x1) y - 2 = -1/5 (x + 3) y-2 = -1/5x - 3/5 y = -1/5x-3/5+2 y = -1/5x+7/5